OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 zero x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

have u ever heard of the matrix theorem?

HanAkoSolo (jamierox4ev3r):

also…couldn't you solve by elimination? which are you familiar with? @Niall ?

OpenStudy (anonymous):

it think u can do both but ive came to know matrix theorem is a lot easier its just more steps

OpenStudy (mathstudent55):

Since you only want the x value, eliminate the y variable and solve for x. Multiply the first equation by -4. Multiply the second equation by 3. Then add the two equations together. The y variable will be eliminated, and you'll have a single equation with only x that you can solve for x.

OpenStudy (anonymous):

niall what do u prefer?

OpenStudy (anonymous):

No, I haven't heard of the matrix theorem, and yeah, I'm familiar with the elimination method

OpenStudy (anonymous):

I prefer the one thats easier :P

OpenStudy (anonymous):

@mathstudent55 The first step is this, right? *-4: 2x − 3y = 3 -8x + 12y = -12

OpenStudy (mathstudent55):

Why don't you try elimination as I explained above. Can you multiply the first equation by -4 and show what you get?

OpenStudy (mathstudent55):

Yes, correct.

OpenStudy (mathstudent55):

Now multiply the second equation by 3.

OpenStudy (anonymous):

5x − 4y = 4 15x - 12y = 12

OpenStudy (mathstudent55):

Great. Using the new multiplied equations, now write the first equation and the second equation just below it, lining up the variables. Notice that the y variable terms are 12y and -12y which add to zero.

OpenStudy (mathstudent55):

Now add the equations -8x + 12y = -12 (+) 15x - 12y = 12 -------------------------

OpenStudy (anonymous):

So would the final answer be 7?

OpenStudy (mathstudent55):

We're solving it together. Let's continue. Now add the equations together. I already set it up for you above.

OpenStudy (anonymous):

-8 + 15 = 7x +12 + -12 = 0[x] And again, +12 + -12 = 0

OpenStudy (mathstudent55):

-8x + 12y = -12 (+) 15x - 12y = 12 ------------------------- 7x = 0

OpenStudy (mathstudent55):

You need to keep the variables.

OpenStudy (mathstudent55):

Now you have a simple equation in only x, so you can solve it.

OpenStudy (mathstudent55):

7x = 0 Divide both sides by 7: x = 0

OpenStudy (anonymous):

So 0's the value of the x variable?

OpenStudy (mathstudent55):

Correct. In this system, x = 0.

OpenStudy (mathstudent55):

That is one of the choices.

OpenStudy (anonymous):

Ok, thanks a ton!!!! I like it way better when people actually take the time to explain how to find the answer, than when they just give it to you! =D

OpenStudy (anonymous):

Can you explain how to solve this too, please? The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (mathstudent55):

Let's call the number of one-step equations x and the number of two-step equations y.

OpenStudy (mathstudent55):

The total number of one-step and two-step equations is 1070. Can you use x and y to write the above statement as an equation?

OpenStudy (anonymous):

Um x + y = 1070?

OpenStudy (mathstudent55):

Great.

OpenStudy (mathstudent55):

Now we need to use again x and y to translate the other statement into an equation.

OpenStudy (mathstudent55):

"If three times the number of one-step equations minus twice the number of two-step equations is equal to 110" We need to write the above statement as an equation.

OpenStudy (anonymous):

Ok, so: 3x - 2y = 110?

OpenStudy (mathstudent55):

three times the number of one-step equations ---> 3x twice the number of two-step ---> 2y

OpenStudy (mathstudent55):

Great. You're ahead of me.

OpenStudy (mathstudent55):

Now we have a system of two equations in two unknowns. x + y = 1070 3x - 2y = 110

OpenStudy (mathstudent55):

Notice the question is "how many one-step equations auditioned today?" Since we let x represent one-step equations, we only need to solve for x.

OpenStudy (mathstudent55):

Adding the equations together as they are will not help because it will not eliminate any variable. We want to eliminate y since we are only interested in x. If we add y (of the first equation) and -2y (of the second equation) together, we get -y. We want zero y, not -y.

OpenStudy (anonymous):

Ohh I see

OpenStudy (mathstudent55):

All we need to do is to multiply the first equation by 2. This way the y term of the first equation will be 2y, which when added to the -2y of the second equation will give us zero y and will eliminate the y variable.

OpenStudy (mathstudent55):

Now we multiply the first equation by 2 and write the second equation below.

OpenStudy (anonymous):

So 2x + 2y = 2140 -------------- 3x - 2y = 110

OpenStudy (mathstudent55):

x + y = 1070 ---- (times 2) ----> 2x + 2y = 2140 3x - 2y = 110 ---- (just copy) ----> 3x - 2y = 110 (add)-------------- 5x = 2250

OpenStudy (mathstudent55):

Yes, you are correct. Now you add the equations you wrote.

OpenStudy (anonymous):

So then you would divide 2250 by 5? :)

OpenStudy (mathstudent55):

Right. Since we ended up with 5x = 225, we divide both sides by 5 to find x. On the right side, that means divide 2250 by 5.

OpenStudy (anonymous):

x = 450

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Okay! Thank you sooooo much!!! =D I really appreciate the help!

OpenStudy (mathstudent55):

Since x represents the number of one-step equations the answer is "450 one-step equations"

OpenStudy (mathstudent55):

You're welcome.